OpenStudy (kirbykirby):

Using exp(iθ) = cos(θ)+ i sin(θ), show that cos^2(θ) + sin(θ) = 1

OpenStudy (anonymous):

You mean \( \sin^2(\theta)\)

OpenStudy (kirbykirby):

^Yes, sorry!

OpenStudy (anonymous):

It is easy \[ e^{i \theta} = \cos(\theta) + i \sin(\theta)\\ e^{-i \theta} = \cos(\theta) - i \sin(\theta)\\ \]

OpenStudy (anonymous):

Multiply the two equations and you are done

OpenStudy (anonymous):

\[ e^{i \theta} = \cos(\theta) + i \sin(\theta)\\ e^{-i \theta} = \cos(\theta) - i \sin(\theta)\\ 1= e^{i \theta} e^{-i \theta} = \cos^2(\theta) -(i \sin(\theta)^2=\cos^2(\theta)+\sin^2(\theta) \]

OpenStudy (anonymous):

Did you understand it?

OpenStudy (kirbykirby):

ohh I see Oh my yes thank you !! :)

OpenStudy (anonymous):

YW